OpenStudy (anonymous):

find the angle in between the given vectors u=<-9,1> v=<-8,7>

OpenStudy (anonymous):

how do i do this i know you take the dot product and divide by the magnitude of u and v but then what?

OpenStudy (phi):

\[u \cdot v= |u| |v| \cos \theta\]